Ab-Initio Calculation of Magnetic Anisotropy in Transition Metal Compounds: Correlation Effects

Gabriel Kotliar Rutgers University

Department of Physics and Astronomy Serin Physics Laboratory Rutgers University 136 Frelinghuysen Road POBox 849 Piscataway N.J 08855

Progamr Officer: Krystl Hathaway

Subcontractors: None

Productivity: From May 1st 1999 till present I have devoted substantial effort to this project. The research goals have been twofold, formulating the evaluation of the magnetic anisotropy constant in cubic crystals as a response function, in order to Incorporating spin orbit coupling to the existent realistic dynamical mean field programs.

Technology Transfer: codes.

1 journal publications

2 Invited Presentations, Honors etc.

Involved in the cause of strongly correlated electron systems.